OpenStudy (anonymous):

5v-4/10=4/5 (please show work)

OpenStudy (anonymous):

multiply both sides by 10 to cancel the ten in the denominator. so you'll have 5v-4 = 8 then add 4 to both sides, so you will have 5v = 12, then divide both sides by 5 so v = 12/5

OpenStudy (anonymous):

you start by multiplying both sides by 10

OpenStudy (anonymous):

(5v-4)5=4*10 ....25v-20=40 ....25v=40+20 ....25v=60 ....v=60/25 .....v=2.4